OpenStudy (anonymous):

Find the equation (in slope intercept form) of the line passing through the points with the given coordinates: (-5,4), (-1,-3)

OpenStudy (anonymous):

Do you know how to find the slope? Do you understand point slope formula?

OpenStudy (anonymous):

\[change \in y/change \in x\]

OpenStudy (anonymous):

slope is \[\frac{4+3}{-5+1}=-\frac{7}{4}\] now use almighty point slope formula \[y-y_1=m(x-x_1)\] with \[y_1=4, x_1=-5, m=-\frac{7}{4}\] to get \[y-4=-\frac{7}{4}(x+5)\] \[y-4=-\frac{7}{4}x-\frac{35}{4}\] \[y=\frac{7}{4}x-\frac{19}{4}\]

OpenStudy (anonymous):

unless my algebra is wrong which is always possible